# arXiv:2601.18474v3[math.NT] 27 Feb 2026

### v3

@@ 1. Introduction
$\begin{array}[]{c}\Gamma^{\left(1\right)}\left(1\right)=-\gamma,\\ \Gamma^{\left(2\right)}\left(1\right)=\gamma^{2}+\zeta\left(2\right),\\ \Gamma^{\left(3\right)}\left(1\right)=-\gamma^{3}-3\zeta\left(2\right)\gamma-2\zeta\left(3\right),\\ \Gamma^{\left(4\right)}\left(1\right)=\gamma^{4}+6\zeta\left(2\right)\gamma^{2}+8\zeta\left(3\right)\gamma+\dfrac{27}{2}\zeta\left(4\right),\\ \Gamma^{\left(5\right)}\left(1\right)=-\gamma^{5}-10\zeta\left(2\right)\gamma^{3}-20\zeta\left(3\right)\gamma^{2}-\dfrac{135}{2}\zeta\left(4\right)\gamma-20\zeta\left(2\right)\zeta\left(3\right)-24\zeta\left(5\right),\\$ $\vdots\end{array}$ (2)
@@
$\begin{array}[]{c}\Gamma^{\left(1\right)}\left(\tfrac{1}{2}\right)=-\sqrt{\pi}g,\\ \Gamma^{\left(2\right)}\left(\tfrac{1}{2}\right)=\sqrt{\pi}\left(g^{2}+3\zeta\left(2\right)\right),\\ \Gamma^{\left(3\right)}\left(\tfrac{1}{2}\right)=-\sqrt{\pi}\left(g^{3}+9\zeta\left(2\right)g+14\zeta\left(3\right)\right),\\ \Gamma^{\left(4\right)}\left(\tfrac{1}{2}\right)=\sqrt{\pi}\left(g^{4}+18\zeta\left(2\right)g^{2}+56\zeta\left(3\right)g+\dfrac{315}{2}\zeta\left(4\right)\right),\\ \Gamma^{\left(5\right)}\left(\tfrac{1}{2}\right)=-\sqrt{\pi}\left(g^{5}+30\zeta\left(2\right)g^{3}+140\zeta\left(3\right)g^{2}+\dfrac{1575}{2}\zeta\left(4\right)g+420\zeta\left(2\right)\zeta\left(3\right)+744\zeta\left(5\right)\right),\\$ $\vdots\end{array}$ (3)
@@ 2.2. Lower Bound for Transcendental Densities
Although the indicated lower bound ($\beta\left(N\right)$) is unaffected by $q$, this does not imply that $\delta_{q}\left(N\right)$ itself is constant over $q\in\frac{1}{2}\mathbb{Z}\setminus\mathbb{Z}_{\leq 0}$. After all, $\beta\left(N\right)\asymp N^{-1/2}$ converges to 0 in the limit, thus providing a rather weak (and not very informative) lower bound. At the same time, it is worth noting that this asymptotic bound is actually stronger (i.e., larger in magnitude) than the posited asymptotic lower bound of $\widetilde{\beta}\left(N\right)\asymp\ln\left(N\right)/N$ given by Theorem 1(iv) of Powers [10]. [^6] It also may be of interest to note that $\beta\left(N\right)$ is larger in magnitude than the asymptotic lower bound for the density of irrational odd-$n$ $\zeta\left(n\right)$ among $n\in\left\{3,5,\ldots,N\right\}$ given by Fischler, Sprang, and Zudilin [9]: $\beta_{\zeta}\left(N\right)\asymp\dfrac{2^{\left(1-\varepsilon\right)\ln\left(N\right)/\ln\left(\ln\left(N\right)\right)}}{\left(N-1\right)/2}\asymp N^{\left(1-\varepsilon\right)\ln\left(2\right)/\ln\left(\ln\left(N\right)\right)-1},$ for arbitrarily small $\varepsilon>0$. More importantly, the new bound is unconditional, whereas the earlier one is conditional on an unproved (albeit computationally supported) conjecture.

### v4

# “Infinitely Often” Transcendence of Gamma-Function Derivatives

@@ 1. Introduction
$\begin{array}[]{c}\Gamma^{\left(1\right)}\left(1\right)=-\gamma,\\ \Gamma^{\left(2\right)}\left(1\right)=\gamma^{2}+\zeta\left(2\right),\\ \Gamma^{\left(3\right)}\left(1\right)=-\gamma^{3}-3\zeta\left(2\right)\gamma-2\zeta\left(3\right),\\ \Gamma^{\left(4\right)}\left(1\right)=\gamma^{4}+6\zeta\left(2\right)\gamma^{2}+8\zeta\left(3\right)\gamma+\dfrac{27}{2}\zeta\left(4\right),\\ \Gamma^{\left(5\right)}\left(1\right)=-\gamma^{5}-10\zeta\left(2\right)\gamma^{3}-20\zeta\left(3\right)\gamma^{2}-\dfrac{135}{2}\zeta\left(4\right)\gamma-20\zeta\left(2\right)\zeta\left(3\right)-24\zeta\left(5\right),\\$ $\cdots\end{array}$ (2)
@@
$\begin{array}[]{c}\Gamma^{\left(1\right)}\left(\tfrac{1}{2}\right)=-\sqrt{\pi}g,\\ \Gamma^{\left(2\right)}\left(\tfrac{1}{2}\right)=\sqrt{\pi}\left(g^{2}+3\zeta\left(2\right)\right),\\ \Gamma^{\left(3\right)}\left(\tfrac{1}{2}\right)=-\sqrt{\pi}\left(g^{3}+9\zeta\left(2\right)g+14\zeta\left(3\right)\right),\\ \Gamma^{\left(4\right)}\left(\tfrac{1}{2}\right)=\sqrt{\pi}\left(g^{4}+18\zeta\left(2\right)g^{2}+56\zeta\left(3\right)g+\dfrac{315}{2}\zeta\left(4\right)\right),\\ \Gamma^{\left(5\right)}\left(\tfrac{1}{2}\right)=-\sqrt{\pi}\left(g^{5}+30\zeta\left(2\right)g^{3}+140\zeta\left(3\right)g^{2}+\dfrac{1575}{2}\zeta\left(4\right)g+420\zeta\left(2\right)\zeta\left(3\right)+744\zeta\left(5\right)\right),\\ \cdots,\end{array}$ (3)
@@ 2.2. Lower Bound for Transcendental Densities
Although the indicated lower bound ($\beta\left(N\right)$) is unaffected by $q$, this does not imply that $\delta_{q}\left(N\right)$ itself is constant over $q\in\frac{1}{2}\mathbb{Z}\setminus\mathbb{Z}_{\leq 0}$. After all, $\beta\left(N\right)\asymp N^{-1/2}$ converges to 0 in the limit, thus providing a rather weak (and not very informative) lower bound. [^6] It may be of interest to note that $\beta\left(N\right)$ is larger in magnitude than the asymptotic lower bound for the density of irrational $\zeta\left(n\right)$ among $n\in\left\{3,5,\ldots,N\right\}$ given by Fischler, Sprang, and Zudilin [9]: $\beta_{\zeta}\left(N\right)\asymp\dfrac{2^{\left(1-\varepsilon\right)\ln\left(N\right)/\ln\left(\ln\left(N\right)\right)}}{\left(N-1\right)/2}\asymp N^{\left(1-\varepsilon\right)\ln\left(2\right)/\ln\left(\ln\left(N\right)\right)-1},$ for arbitrarily small $\varepsilon>0$.